OpenStudy (anonymous):

An investor invested a total of $2,700 in two mutual funds. One fund earned a 4% profit while the other earned a 2% profit. If the investor's total profit was $94. How much was invested in each mutual fund?

OpenStudy (anonymous):

You don't know at first how much was in each fund. However, you do know that the total amount was $2700. So you could say that the amount in the 4% fund was "x" which then means the amount in the other 2% fund was everything leftover... 2700-x

OpenStudy (anonymous):

Thankfully, you also know the total profit was $94. This is enough to set up an equation using the profit and the expressions I showed in the last post. Profit = (amount in fund 1 * fund 1 profit %) + (amount in fund 2 * fund 2 profit %) 94 = (x * 4%) + ( (2700 - x) * 2%)

OpenStudy (anonymous):

That equation would be easier to solve for x if you used decimals rather than percentages... 94 = (x * 4%) + ( (2700 - x) * 2%) 94 = 0.04x + 0.02(2700 - x) Then simplify and solve for x. That is the amount in the 4% fund, and 2700-x is the amount in the 2% fund.

OpenStudy (anonymous):

Can you please explain how to do that. I need to know the amount of each. I am not just looking for the answer. I need to know how to do the problem. Thank you.